# Automatically Analyzing Performance Issues in Android Apps: How Far Are We?

### v1

@@ 3.1. Data Collection
Fig. 3 shows our three-step pipeline to obtain real-world performance issues, discussions, and solutions. Briefly, we first crawl user reviews from the Google Play Store, question threads from Stack Overflow, and issue-commit pairs from GitHub. Next, we filtered collected information by keywords obtained from RQ1. Lastly, we manually verified real-world performance issues, discussions, and their common patterns.
@@ 3.1.1. Data Crawling
To understand real-world performance issues, we collect data from three popular platforms in the Android development ecosystem as below.
@@
[Google Play - User Reviews.] The Google Play Store is the biggest Android app store in the world. It constitutes an essential component of the Android ecosystem, facilitating the distribution and discovery of apps for millions of users worldwide. Users frequently complain about the performance issues of the Android apps by leaving reviews in the Google Play Store, providing useful information for researchers (Liu2014CharacterizingAD, ). We collected user reviews from the apps on Google Play based on the AppBrain popularity ranking [^2] AppBrain website https://www.appbrain.com/stats/google-play-rankings. To hold the representativeness, we collected the top 200 apps in five monetary categories (top free, top paid, top grossing, top new free, top new paid) across 30 countries, totaling 7,681 unique apps. We used the Google Play scraper library333google-play-scraper: https://pypi.org/project/google-play-scraper/ and successfully collected 909,430 user reviews from these apps. Please note that we excluded user reviews which only have a rating score (1-5) without comments. Intuitively, performance issues are embedded in the negative user reviews, so we used the sentiment analysis model SiEBERT (hartmann2023, ) to identify negative reviews. This model is a fine-tuned checkpoint of RoBERTa (Liu2019RoBERTaAR, ), with fine-tuning datasets from diverse text sources such as reviews and tweets. As a result, we roughly obtained 60,684 negative user reviews for further process.
@@
[Stack Overflow - Question Threads.] Stack Overflow is a popular Q&A website for programmers and developers. It plays a crucial role in the Android ecosystem by providing a platform for knowledge sharing, problem-solving, and community support (Wu2022RetrievingAK, ; Huang2018APIMR, ). Thus, we explored the performance issues developers consider and discuss in the development process. Each question threads on Stack Overflow normally reflects one real-world programming challenges, and the follow-up discussion in the thread provide additional thoughts, insights, and potential solutions. To understand developers’ opinions on Android performance issues, we crawled all question threads labeled with “android” and containing accepted answers, resulting in 448,977 questions. We further narrowed down to 360,372 threads containing the code tag for further refinement.
@@
[Github - Issues & Commits.] GitHub is a critical platform in the Android ecosystem. Issues and commits are essential tools on GitHub that help manage the development process in Android development projects. Issues provide a structured way to handle reported performance issues, while commits ensure a clear and traceable history of code changes as solutions. Therefore, issues and commits can provide valuable insights about how developers actually solve those issues in the maintenance process. We used F-Droid (FDroid, ), an open-source Android app repository, to access each app’s corresponding GitHub repository. Here, we chose F-Droid because it is way more easier to inspect code patterns in open-source apps compared to closed-source ones. To gather representative Android apps, we crawled 17 categories of open-source Android apps and their GitHub links from F-Droid. Based on these links, we successfully identify 4,228 Android apps and their GitHub repositories, containing 35,278 issues and 550,973 commits.
@@ 3.1.2. Keyword Filtering
As we focus on Android performance issues, we adapted a keyword-based filtering to select relevant data only. Alongside of the systematical literature review of 48 academic research papers in Section 2.1.4, we also snowballed a list of keywords about Android performance issues. The keyword list includes terms like “energy waste”, “response time”, “responsive”, and so on. As a previously unknown performance issue might be caused by an existing factor and vice versa, keywords from existing literature do not restrict us from discovering new performance issues and new factors. Using these keywords, we filtered 60,684 user reviews to 132, 360,372 Stack Overflow discussions to 773, 35,278 GitHub issues down to 297, and 550,973 GitHub commits to 645.
@@
After filtering data by performance-related keywords, we notice some data entries are false positive, meaning they contained keywords but were not actually related to performance issues or their root causes. Therefore, we further conducted a manual verification to ensure the data quality and relevance.
User Reviews. We removed those unrelated to performance issues, such as criticisms of developers. For example, “The app works well, but only supports SMBv1, which has been deprecated and known to be very insecure. There have been numerous complaints about this but the author has been stubbornly unresponsive about it…”
Question Threads. For Stack Overflow, we excluded question threads that are wrongly labelled on the code tag and those low quality threads that their negative votes outnumber positive votes.
GitHub - Issues. We first checked and removed issues from inactive (dead) projects, such as those where developers do not respond or have explicitly stated that the project is no longer maintained. Additionally, we excluded issues that do not involve performance issues, such as “Show me … ports mac address host name mDNS response time, perhaps option to ping again and option to traceroute.
GitHub - Commits. We manually checked all remaining commits, and removed those without clear commit messages. Additionally, we excluded commits with multiple changes, such as fixing various bugs or features, where performance was only a minor part, making it difficult to identify the root causes.
Above all, we obtained 104 user reviews, 332 stack overflow discussions, 78 GitHub issues, and 289 GitHub commits. Two of authors performed the aforementioned manual checking individually, for any disagreement, they discussed and reached the same answer. For persisting disagreement, a senior author will then involve and decide. We utilized Cohen’s kappa (Landis1977AnAO, ) to assess inter-rater agreement. The results indicate a Cohen’s kappa of 0.953 for user reviews, 0.897 for commits, 0.923 for GitHub issues, and 0.885 for Stack Overflow posts, all of which signify substantial agreement. We then manually identified the Android performance issues in these data and present our findings in the next section.
@@
Figure 4. Distribution of performance issues and factors in user reviews, stack overflow, GitHub issues, and GitHub commits. Other Factors: Polling, Syntax, Hasmap, Cookies, Override, Thread Pool, Cache, Copy Text, Background Thread, HTTP Request.
@@ 3.2. Result Analysis
Fig. 4 illustrates the distribution of performance issues and factors across the three data sources. Overall, we identified six categories of performance issues (including one previously unseen) and 35 factors (including 19 previously unseen) from the three real-world data sources. Next, we analyzed our findings from the perspectives of User Reviews, Stack Overflow, and GitHub Issues & Commits.
3.2.1. Performance Issues and Factors in Real-world
We first identified performance issues in real-world apps and highlight the gaps between these issues and those discussed in research communities.
@@ 3.2.1. Performance Issues and Factors in Real-world
[Google Play - User Reviews.] User reviews typically focus on how performance issues negatively impact their experience without delving into the specific causes. Some user reviews mention more than one performance issue, so the percentages add up to more than 100%. We totally identified six performance issues, five of which were known from the literature review, while one performance issue was previously unseen. Responsiveness issues, such as slow responsiveness or UI freezes, have the most direct impact on user experience, appearing 79 times (75.96%). For example, one comment says: “… the software appears to be sluggish in map updating and does not properly locate my vehicle… I am hopeful these issues will be resolved in the future but if not then I will stop discontinue my premium subscription service and find another solution.” This indicates that unresolved performance issues can lead to users canceling subscriptions, causing companies to lose market share. Energy Consumption is also critical, as inefficient app power consumption can quickly drain batteries. These issues appeared 25 times (24.04%). E.g., one representative comment states: “severe battery drain on s10, causing 1% loss every 3 minutes. drain immediately disappeared after uninstall, back to 1% loss every 8-9 minutes.” This shows that performance issues can lead to users uninstalling the app. Additionally, a performance issue not covered in existing literature is Data Consumption. E.g., one comment mentions: “.. Still cant even disable it, much less delete it…And worst of all, it still runs, using up data, memory and running down battery. Shame on You XXX…”
[Stack Overflow - Question Threads.] From the 332 question threads, we identified five categories of performance issues, all of which are already mentioned in the literature review (Section 2). Developers most frequently discussed Memory Consumption (80.72%), which is mainly caused by Image factor. However, the prevalence of Memory Consumption only ranks forth in existing research papers according to Fig. 2. Additionally, while Responsiveness is the most concerning issue for end users, it’s the second most frequent topic in developer discussions. This discrepancy indicates that the performance issues developers focus on, those users are concerned about, and the ones researchers prioritize, do not align. Moreover, we identified a total of 30 factors contributing to these performance issues, and only 16 factors appeared in the literature review. Among the 14 newly discovered factors, the most frequent one is File Operation, which typically involves inefficiently downloading, uploading, or saving large files, leading to memory shortages. The next most frequent previously unseen factor is Data Type, which often lead to Memory Consumption and Responsiveness issues.
[Github - Issues & Commits.] Among the 289 commits, we identified four categories of performance issues, all of which are already mentioned in the literature review (Section 2). The most frequent performance issue is Memory Consumption (84.43%), followed by Responsiveness (11.07%). This distribution closely resembles that found in Stack Overflow discussions, indicating that developers are most concerned with memory consumption. In addition, we identified 22 factors, 11 of which are previously unseen. For example, the factor Parameter can lead to data loss or memory consumption. In Commit  [^4] https://github.com/ente-io/ente/commit/0165b7569b4acee1699de59c7c6ec1a8b91a7b4f, the change was to remove the fileUrl from console.log parameters to prevent a memory leak.
Regarding GitHub issues, we identified four performance issues and nine factors (including three unseen factors). The most frequently discussed issue is Data Loss (51.28%), followed by Memory Consumption (38.46%) and Responsiveness (8.97%). Interestingly, Data Loss is the most common issue in GitHub Issues, while Memory Consumption is the most frequently addressed in commits. This discrepancy indicates that the issues reported in GitHub issues are not always the ones developers prioritize for fixing.
Answers to RQ2.1 We identified an unseen performance issue related to Data Consumption and 19 previously unrecognized factors such as Data Type, File Operation, and Object Creation, highlighting gaps in research coverage. The primary concerns of researchers, developers, and users are not aligned: researchers mainly focus on energy consumption, developers prioritize memory consumption, and users are most concerned with responsiveness.

### v2

@@ 2.1. Data Collection
Following common practices in analyzing large-scale online discussions (Guo2013CharacterizingAD, ; Liu2014CharacterizingAD, ; Rua2023ALE, ; Banerjee2018EnergyPatchRR, ), we collected data from three popular crowd-sourced platforms: Google Play, Stack Overflow, and GitHub. Indeed, Google Play provides user feedback to assess the impact on end-users, Stack Overflow offers discussions that reveal common issues and solutions, and GitHub contains issues and commits histories that reflect the challenges developers encounter. Fig. 1 shows our three-step pipeline to perform the real-world exploration. Briefly, we first crawl user reviews from the Google Play Store, question threads from Stack Overflow, and issue-commit pairs from GitHub. Then, we filtered this data using performance-related keywords. Lastly, we manually checked real-world performance issues, root causes, and their common code patterns.
@@ 2.1.1. Data Crawling
Here, we elaborate the details of data crawling as follows.
@@
[Google Play - User Reviews.] To collect user reviews, we focus on those from the Google Play Store, the largest Android app marketplace globally. This platform plays a crucial role in the Android ecosystem, enabling app distribution and discovery for millions of users worldwide. Users often report performance issues through their reviews on the Google Play Store, providing valuable insights for researchers (Liu2014CharacterizingAD, ). We collected user reviews from the apps on Google Play based on the AppBrain popularity ranking. To ensure representativeness, we collected the top 200 apps in five monetary categories (top free, top paid, top grossing, top new free, top new paid) across 30 countries, totaling 7,681 unique apps. We used the Google Play scraper library111google-play-scraper: https://pypi.org/project/google-play-scraper/ and successfully collected 909,430 user reviews from these apps. Please note that we excluded user reviews that only have a rating score (1-5) without comments. Intuitively, performance issues are embedded in the negative user reviews, so we used the sentiment analysis model SiEBERT (hartmann2023, ) to identify negative reviews. This model is a fine-tuned checkpoint of RoBERTa (Liu2019RoBERTaAR, ), with fine-tuning datasets from diverse text sources such as reviews and tweets. As a result, we obtained 60,684 negative user reviews for further processing.
@@
[Stack Overflow - Question Threads.] We also collected performance-related question threads from Stack Overflow, a widely recognized Q&A platform for programmers and developers, which plays a crucial role in the Android ecosystem by facilitating knowledge sharing, problem-solving, and community support (Wu2022RetrievingAK, ; Huang2018APIMR, ). Specifically, we explored the performance issues developers posted and discussed in the development process. Each question thread on Stack Overflow typically reflects a real-world programming challenge, with the follow-up discussions offering additional insights, perspectives, and potential solutions. To understand developers’ perspectives on Android performance issues, we crawled all question threads labeled with “android” and containing accepted answers with code, resulting in 360,372 questions.
@@
[Github - Issues & Commits.] Additionally, we collected performance issues and commits for open-source apps from GitHub, the most widely used platform for open-source development. GitHub issues provide structured information of reported performance issues, while commits create a clear and traceable history of code fixes made as solutions. Consequently, both issues and commits offer valuable insights into how developers effectively resolve these issues during the maintenance process. To access each app’s corresponding GitHub repository, we utilized F-Droid (FDroid, ), an open-source Android app repository. We chose F-Droid because it facilitates easier inspection of code patterns in open-source applications compared to closed-source ones. To gather a representative sample of Android apps, we crawled 17 categories of open-source Android apps along with their GitHub links from F-Droid. As a result, we successfully identified 4,228 Android apps and their GitHub repositories, which contained 35,278 issues and 550,973 commits.
@@ 2.1.2. Keyword Filtering
As we focus on Android performance issues, we implemented a keyword-based filtering approach to select relevant data. To compile an accurate keyword list, we conducted a Google search for “Android performance”, reviewing the top 10 most relevant pages, which included 100 websites. Additionally, we examined several foundational papers on Android performance issues (Hort2021ASO, ; Liu2014CharacterizingAD, ; Habchi2019TheRO, ; Rua2023ALE, ; Das2020CharacterizingTE, ). From these sources, we obtained a list of 87 keywords about Android performance issues. The keyword list includes terms like “UI freeze”, “battery drain”, “memory leak”, “cpu utilization”, “app size” and so on. As a previously unknown performance issue might be caused by an existing factor and vice versa, existing keywords do not restrict us from discovering new performance issues and new factors. Using these keywords, we filtered 60,684 user reviews to 165, 360,372 Stack Overflow discussions to 893, 35,278 GitHub issues down to 317, and 550,973 GitHub commits to 780.
@@
After filtering the data using performance-related keywords, we noticed some data entries were false positive, meaning they contained keywords but were not actually related to performance issues. Therefore, we further conducted a manual verification to ensure the data quality and relevance.
User Reviews. We removed user reviews that were unrelated to performance issues, such as criticisms of developers. For example, one review stated: “The app works well, but only supports SMBv1, which has been deprecated and known to be very insecure. There have been numerous complaints about this but the author has been stubbornly unresponsive about it…”
Question Threads. For Stack Overflow, we excluded question threads that are wrongly labeled on the code tag as well as low-quality threads where the number of negative votes exceeded the number of positive votes.
GitHub - Issues. We first checked for and removed issues from inactive (dead) projects, such as those where developers do not respond or have explicitly stated that the project is no longer maintained. Additionally, we excluded issues that do not involve performance issues, such as “Show me … ports mac address host name mDNS response time, perhaps option to ping again and option to traceroute.
GitHub - Commits. We manually checked all remaining commits and removed those that lacked clear messages. Additionally, we excluded commits with multiple changes, such as fixes for various bugs or features where performance was a minor aspect, as this made identifying the root cause difficult.
Ultimately, we obtained 114 user reviews, 336 stack overflow discussions, 47 GitHub issues, and 295 GitHub commits. Two of the authors performed the aforementioned manual checking individually, for any disagreement, they discussed their findings until reaching a consensus. For any persisting disagreement, a senior author made the final decision. We utilized Cohen’s kappa (Landis1977AnAO, ) to assess inter-rater agreement, yielding scores of 0.944 for user reviews, 0.885 for GitHub issues, 0.895 for GitHub commits, and 0.871 for Stack Overflow posts, all of which signify substantial agreement. We then manually identified the Android performance issues within this data and present our findings in the following section.
@@ 2.2. Result Analysis
To gain a thorough understanding of prevalent performance issues in real-world apps, we begin by analyzing the distribution of performance issues encountered by users and developers (Section 2.2.1). Next, we categorize the underlying root causes contributing to these issues (Section 2.2.2). Finally, we identify code patterns commonly associated with these performance issues (Section 2.2.3).
2.2.1. Performance Issues in Real-world
We identified seven key performance issues that are of concern to users and frequently discussed by developers in real-world scenarios. Fig. 2 shows the distribution of these issues across different platforms.
Figure 2. Distribution of performance issues across various sources.
@@ 2.2.1. Performance Issues in Real-world
[Google Play - User Reviews.] User reviews typically focus on how performance issues negatively impact their experience without delving into the specific causes. As shown in the User Review pie chart in Fig. 2, users are mainly concerned about responsiveness, energy consumption, storage consumption, memory consumption, and internet data usage. Responsiveness issues, such as slow responsiveness or UI freezes, have the most direct impact on user experience, appearing in 62.3% of all reported reviews. This is not surprising as responsiveness is the most noticeable aspect of performance that users can directly perceive. For example, one comment says: “… the software appears to be sluggish in map updating and does not properly locate my vehicle… I am hopeful these issues will be resolved in the future but if not then I will stop discontinue my premium subscription service and find another solution.” This indicates that unresolved performance issues can lead to users canceling subscriptions, causing companies to lose market share.
Energy Consumption is also critical, as inefficient app power consumption can quickly drain batteries. These issues appeared in 21.93% of all reported reviews. E.g., one representative comment states: “severe battery drain on s10, causing 1% loss every 3 minutes. drain immediately disappeared after uninstall, back to 1% loss every 8-9 minutes.” This shows that performance issues can lead to users uninstalling the app. Storage Consumption is often complained by users that the size of the app is too large, or the cache space consumed is too large when using the app. These issues appeared in 9.65% of all reported reviews. As mobile devices have limited storage, excessive storage consumption can lead to problems like a lack of available space for other apps or files, and potential app crashes. E.g., one comment mentions: “..Every time a full size version of a Google app gets updated it eats up even more internal storage space… if there’s not enough storage space in your internal storage it won’t install certain updates or new apps from Google Play store, you have to uninstall updates or apps before reinstalling updates and apps…” Finally, Internet Data Usage issues are reported by users when apps consume excessive data, especially when they aren’t using Wi-Fi, making them more sensitive to data usage. For example, one user mentioned: “…Not sure why, but my data usage has skyrocketed. I’m on a limited plan, now I have to buy extra data every month…”
[Stack Overflow - Question Threads.] From the 336 question threads, we identified six categories of performance issues, as illustrated in the Stack Overflow pie chart in Fig. 2. Interestingly, two performance issues (i.e., CPU usage and GPU usage) are not mentioned in user reviews. Additionally, while responsiveness is the most concerning issue for end users, it’s the second most frequent topic in developer discussions. This discrepancy indicates that the performance issues developers focus on and those users are concerned about do not align. Since user satisfaction directly impacts app market share, developers should prioritize addressing the issues most important to users.
[Github - Issues & Commits.] The GitHub Issues and GitHub Commits pie charts in Fig. 2 show a similar distribution of performance issues across both data sources, with memory consumption, responsiveness, and storage consumption being the most frequently mentioned. Notably, while internet data usage is a niche issue primarily highlighted in user reviews and absent from Stack Overflow and GitHub commits, it also appears in GitHub issues, providing insights into its underlying causes.
In summary, the five performance issues that users mention in reviews are also discussed by developers on GitHub and Stack Overflow. Developers further address two additional issues, CPU usage and GPU usage, which are rarely mentioned by users, likely because they are technical concerns not directly perceived in user experience.
Answers to RQ1.1 We identified seven categories of performance issues: responsiveness, energy consumption, storage consumption, memory consumption, internet data usage, CPU usage, and GPU usage. A substantial divergence exists between the focus areas of performance issues for users and developers: while users primarily emphasize concerns around responsiveness, developers’ discussions are predominantly oriented toward memory consumption.
Table 1. Taxonomy of seven categories of performance issues and their contributing factors. ‘R-W.’ indicates real-world sources, while ‘Lit.’ refers to literature sources.
Category Factor R-W. Lit. Problem Example
Memory Consumption Loop Repeatedly reading the same value into the array in an infinite loop
Object Creation Repeatedly creating new instances instead of reusing existing ones
Massive Data Outputting thumbnail URLs containing a large amount of information to the log
Database Operation Each database query loads new data into memory without clearing the previous data
Activity An anonymous inner class strongly references the Activity, preventing it from being garbage collected
File Handling Loading an entire large file into memory at once, rather than processing it in chunks
Fragment The previous fragment state isn’t properly cleared when switching between fragments
Dynamic UI Loading Recreating the view and bitmap for each list item during scrolling, without effectively reusing the views
Background Thread Failing to stop the refresh when the app enters the background causes the UI component to remain active
Singleton A singleton holds a reference to a UI component after its Activity is destroyed, preventing garbage collection
Override Failing to call super.onDestroy() after overriding a method prevents proper release of system resources
Configuration Gradle JVM heap size is insufficient during project builds, causing OOM errors
Service Failing to destroy a dead service  (Song2019ServDroidDS, )
Resource Improper releasing of resources (e.g., Sensor, Camera)  (Liu2014GreenDroidAD, ; Liu2013WhereHM, ; Li2019DetectingAD, ; Behrouz2016EnergyawareTM, ; Li2022CombattingEI, ; Liu2016FixingRL, ; Couto2020EnergyRF, ; Jabbarvand2020AutomatedCO, ; Li2020EnergyDxDE, ; Lu2022DetectingRU, ; Banerjee2018EnergyPatchRR, ; Wu2016LightWeightIA, ; Jabbarvand2019SearchBasedET, ; Banerjee2014DetectingEB, ; Yan2013SystematicTF, ; Liu2019DroidLeaksAC, )
Third-Party Library The API usage in the new version of the third-party library caused excessive memory usage (Rua2023ALE, )
UI Rendering Reinflating views on each “getView” call without reusing them increases memory usage (Das2020CharacterizingTE, )
Data Type Declaring a variable as static holds a strong reference to the previous task, preventing garbage collection (Das2020CharacterizingTE, )
Obsolete Task Retaining obsolete parameters after updating view layouts  (Kang2016DiagDroidAP, ; Couto2020EnergyRF, )
Responsiveness Service Failing to move the notification to the foreground in time during service start, causing an ANR
Inefficient Method Selecting a less efficient method introduces unnecessary delays in processing UI actions
Background Thread Rapidly creating numerous threads in the background to perform I/O on the same data
Cache Large cache causes frequent I/O operations, leading to lag
Text Handling Repeatedly calling EditText.setText() to highlight all search results at once leads to unnecessary UI updates
Resource Insufficient thread resource allocation causes slow response during simultaneous operations (Bhatt2020AutomatedRO, )
UI Rendering Improperly stacking and layering in layouts can result in pixels being redrawn multiple times in the same frame, with only the last rendering visible (Gao2017EveryPC, ; Lin2024AgingOG, ; Liu2014CharacterizingAD, ; Habchi2019TheRO, ; Vsquez2015HowDD, ; Rua2023ALE, ; Das2020CharacterizingTE, ; Janssen2022OnTI, )
Loop An infinite while loop in the button’s onClick() event makes the app unresponsive after clicking (Lyu2018RemoveRF, )
Main Thread Performing heavy operations in the main thread  (Liu2014CharacterizingAD, ; Brocanelli2018HangDR, ; Lin2014RetrofittingCF, ; Gmez2016MiningTR, ; Vsquez2015HowDD, ; Das2020CharacterizingTE, )
Activity Loading new resources during activity restarts may freeze the UI  (Guo2022DetectingAF, ; Chen2023TransparentRC, ; Lu2022DetectingRU, )
View Allocating new objects along with draw operations (Habchi2019TheRO, ; Couto2020EnergyRF, ; Das2020CharacterizingTE, )
Obsolete Task Downloading obsolete content occurs when switching activities  (Kang2016DiagDroidAP, ; Couto2020EnergyRF, )
Advertisements Loading ads while maintaining responsiveness to user actions (Petalotis2024AnES, )
Message Backlogging rapidly incoming messages in Message Handler’s single-threaded (Kang2016DiagDroidAP, )
Storage Consumption File Handling Temp files from URIs aren’t cleared after upload
Redundant Code Similar functional pages and activities that could be reused, or unnecessary code that only occupies storage
Configuration Including unnecessary third-party library dependencies in the Gradle file
Image PNG images compress less effectively than JPG, resulting in larger file sizes
Energy Consumption Resource Updating unnecessary sensor data in the background  (Liu2014GreenDroidAD, ; Liu2013WhereHM, ; Li2019DetectingAD, ; Behrouz2016EnergyawareTM, ; Li2022CombattingEI, ; Liu2016FixingRL, ; Couto2020EnergyRF, ; Jabbarvand2020AutomatedCO, ; Li2020EnergyDxDE, ; Guo2013CharacterizingAD, ; Lu2022DetectingRU, ; Banerjee2018EnergyPatchRR, ; Wu2016LightWeightIA, ; Jabbarvand2019SearchBasedET, ; Banerjee2014DetectingEB, ; Yan2013SystematicTF, ; Jabbarvand2017DroidAE, ; Nikzad2014APEAA, ; Wu2016Relda2AE, ; Vsquez2015HowDD, ; Bhatt2020AutomatedRO, ; Rua2023ALE, ; Liu2019DroidLeaksAC, )
Database Operation Frequently triggering database queries for auto-complete (Li2019DetectingAD, ; Lyu2018RemoveRF, ; Li2022CombattingEI, ; Lyu2017AnES, )
Loop Executing energy-intensive behaviors in loop  (Banerjee2014DetectingEB, ; Behrouz2016EnergyawareTM, ; Lyu2018RemoveRF, ; Fratantonio2015CLAPPCL, ; Lyu2017AnES, ; Jabbarvand2017DroidAE, ; Li2014AnES, ; Palomba2019OnTI, )
HTTP Request Dividing a large packet into small pieces and transmitting data frequently (Li2016AutomatedEO, ; Li2014AnES, )
Image Decoding identical image multiple times (Song2021IMGDroidDI, ; Li2019CharacterizingAD, ; Carette2017InvestigatingTE, ; Kang2016ExperienceRD, )
Wakelock Forgetting to release wakelocks after usage  (Liu2014GreenDroidAD, ; Pathak2012WhatIK, ; Liu2016UnderstandingAD, ; Li2022CombattingEI, ; Couto2020EnergyRF, ; Li2020EnergyDxDE, ; Behrouz2016EnergyawareTM, ; Banerjee2018EnergyPatchRR, ; Banerjee2014DetectingEB, ; Pathak2012WhereIT, ; Jabbarvand2017DroidAE, ; Vsquez2015HowDD, ; Das2020CharacterizingTE, )
UI Rendering 1) Views are rendered but not visible to users because they are off-screen or set to invisible  (Gao2017EveryPC, ; Lin2024AgingOG, ; Liu2014CharacterizingAD, ; Habchi2019TheRO, ; Janssen2022OnTI, )
Inefficient Method Calling unnecessary methods (Carette2017InvestigatingTE, ; Couto2020EnergyRF, ; Habchi2019TheRO, ; Palomba2019OnTI, )
Display Screen The OLED display consumes more energy when displaying light colors than dark ones (Li2014MakingWA, ; LinaresVsquez2018MultiObjectiveOO, ; Vsquez2015OptimizingEC, ; Wan2015DetectingDE, )
Program Language Developing web applications using JavaScript consumes more energy than WebAssembly (Hasselt2022ComparingTE, ; Oliveira2017ASO, )
Advertisements Frequent advertisement and analytics (A&A) service requests lead to energy waste (Cito2016BatteryawareTI, ; Petalotis2024AnES, )
Machine Learning The inefficient machine learning algorithm on mobile devices is overly complex (McIntosh2018WhatCA, )
View Drawing all items in a ListView without reusing existing data  (Habchi2019TheRO, ; Couto2020EnergyRF, )
CPU Usage Multi-Threads The music player mishandles metadata extraction, spawning thousands of threads, overloading the CPU.
Loop An infinite loop repeatedly calls methods to handle images
Image Glide consumes significant CPU resources to sequentially display images when playing GIFs
UI Rendering Unnecessarily using transparency to render UI components (Lee2019ImprovingEE, )
GPU Usage Third-Party Library Third-party library’s bug causes continuous redrawing when using GPU to render UI components
UI Rendering Spin animation has unnecessary frequent refreshes, forcing GPU rendering on every frame (Lee2019ImprovingEE, )
Internet Data Usage Resource Downloading unwanted content without user permission, wasting internet resources
Image Automatically loading animated GIFs without Wi-Fi enabled
Advertisements Automatically downloading embedded advertising services in the background
2.2.2. Root Causes of Performance Issues
To address performance issues, it is crucial to investigate the underlying root causes that lead to these issues. Based on our real-world analysis, we manually summarized the root causes (i.e., contributing factors) for each performance category, as indicated by the checkmarks (✓) in the R-W. column in Table 1.
Here, we remind readers that during the process of manual summarization, we found that the descriptions of performance issues are chaotic. The absence of a universally accepted taxonomy results in many discussions targeting the same performance problems but using different descriptive terms (Hort2021ASO, ; Liu2014CharacterizingAD, ; Habchi2019TheRO, ; Rua2023ALE, ; Das2020CharacterizingTE, ). Additionally, since one factor can lead to multiple performance issues consequences, and one consequence can be caused by multiple factors, it complicates the creation of a clear and comprehensible performance classification system. To address this gap, our work distinguishes between the consequences and factors of performance issues. Specifically, we first classify performance issues by their consequences, then define a series of factors that cause these issues, and finally link the consequences to their corresponding factors. To the best of our knowledge, we are the first to systematically classify all Android performance issues in a hierarchical manner. We now elaborate on the seven major categories, each with the corresponding 49 factors along with an example presented in Table 1.
Memory Consumption. As discussed in Section 2.2.1, memory consumption is the most frequently highlighted performance issue among developers. Our real-world analysis revealed 20 factors that contribute to these memory consumption issues, which are listed under the Memory Consumption category in Table 1. These issues arise from a variety of factors, often resulting from inefficient coding practices that lead to excessive or unnecessary memory use. For example, repeatedly reading the same value into an array within an infinite loop can cause memory to be continuously occupied without being released, ultimately leading to an out-of-memory (OOM) error. Additionally, failing to destroy an AsyncTask when it holds references to active GUI components can prevent the garbage collection of the destroyed GUI, leading to memory leaks.
Responsiveness. Another important aspect of Android performance issues is responsiveness. Poor responsiveness leads to perceptible delays such as image rendering delays, UI freezes, unresponsiveness, etc. In our analysis, we identified 14 key factors contributing to this issue. Many real-world scenarios show that heavy operations associated with UI threads can slow responsiveness, such as decoding large images in the UI threads. Besides, inefficient code logic can also slow down the response time. For example, improper stacking and layering in layouts can cause pixels to be redrawn multiple times within the same frame, slowing down the response time.
Storage Consumption. As the third most concerning performance issue for developers, inefficient storage consumption often leads to apps occupying more storage space than expected, which can limit the device’s ability to install other apps and store data. Through our real-world exploration, we identified four key factors contributing to storage consumption issues. These factors typically arise from inefficient coding practices and poor data management, resulting in excessive or unnecessary storage usage. For instance, creating a large number of similar functional pages and activities that could have been reused, or including redundant code that never gets executed during runtime, wastes storage space. Moreover, adding unnecessary third-party library dependencies to the Gradle file further increases the app’s size after building, leading to larger APK files.
Energy Consumption. As outlined in Section 2.2.1, energy consumption ranks as the second highest concern among users, as apps that drain battery faster than expected can frustrate users, especially given the limited battery resources on mobile devices. However, developers have not prioritized addressing this issue to the same extent. Consequently, in our real-world analysis, we identified only two key factors contributing to energy consumption problems. The root causes generally stem from inefficient resource usage and unnecessary background operations that provide no tangible benefits to users. For instance, updating unneeded sensor data in the background or running background services at unnecessarily high frequencies both result in excessive energy consumption. It is worth noting that, although developers do not prioritize energy efficiency, many research papers have extensively studied this issue, bridging the gap in developer attention. We will further discuss the performance issues targeted by researchers in Section 3.2.1.
CPU Usage. Excessive CPU usage can cause the device to overheat, triggering the thermal management system to automatically lower the processor’s frequency to prevent overheating. This leads to a decline in overall system performance, affecting the normal operation of other apps and background services, and may even result in crashes, while also shortening the lifespan of the hardware and battery. Based on our analysis, we identified four distinct factors contributing to CPU usage issues. These issues typically stem from unnecessary or inefficient operations that waste CPU resources. For instance, using transparency in canvases where it’s not needed forces the browser to perform extra work, increasing CPU load. Similarly, auto-playing GIFs by sequentially displaying images instead of using a static preview image results in unnecessary CPU consumption.
GPU Usage. The GPU is essential for handling graphics-intensive tasks like UI rendering, animations, 3D graphics, and video/image processing, reducing the CPU’s workload. Improper or excessive GPU use can lead to issues like UI freeze, frame drops, device overheating, rapid battery drain, and inefficient system resource allocation. Our real-world analysis identified two key factors contributing to GPU usage issues. These issues typically arise from overly frequent rendering requests or third-party libraries’ bugs. For example, a spin animation with excessively frequent refreshes forces the GPU to render every frame, unnecessarily consuming resources. Besides, a bug in a third-party library could cause an app’s misbehavior such as continuous redrawing of UI components, resulting in excessive GPU usage without providing any tangible benefits.
Internet Data Usage. Network resources on mobile devices are limited, especially when Wi-Fi is unavailable, requiring users to purchase cellular data. Inefficient data usage can lead to unnecessary expenses and create a negative user experience. From our analysis, we identified three key factors contributing to this issue. The root causes often involve the unnecessary consumption of network resources. For example, downloading unwanted content without user consent or automatically downloading embedded advertisements in the background can lead to excessive data usage.
Answers to RQ1.2: Performance issues are inherently complex, often arising from an interplay of multiple contributing factors, where a single issue may stem from various causes, and, conversely, a single factor may amplify multiple issues. We identified 49 distinct factors contributing to the seven categories of performance issues.